# black holes; early universe; general relativity

### v1

@@ 4 Initial spins
If the black hole no hair conjecture holds in astrophysics, stationary black holes in vacuum should be well approximated by Kerr black holes which are characterised only by two parameters, the mass and the angular momentum. As we have already discussed, the mass of PBHs is approximately equal to that within the Hubble horizon at the time of formation. So, what determines the spin of PBHs? If black holes of masses from several to several tens of solar masses are observed, it would be very difficult to distinguish between PBHs and astrophysical black holes. Astrophysical black holes form in the final stage of evolution of massive stars. If we observe an isolated black hole, the only information of its own other than its mass is its spin. So, if the spins of PBHs are expected to be very different from those of astrophysical black holes, the observation of spin can be potentially decisive information to distinguish between the two populations. Here, we focus on the initial spins of PBHs, although the accretion and merger history after their formation may also greatly affect their spins depending on the scenarios De Luca et al. (2020).
@@ 4.1 Spins of primordial black holes formed in radiation domination
It was discussed that the formation process of PBHs in radiation domination is well approximated by spherically symmetric dynamics since Carr (1975) Carr (1975). Recent quantitative studies have revealed that this early argument is basically correct. The key of the quantitative analysis is peak theory.
@@
For PBHs formed in radiation domination, the threshold density perturbation $\delta_{\rm th}$ is of the order of the unity, which is considerably larger than its standard deviation. In other words, PBHs can form only at very rare peaks. Peak theory predicts that there is only very small deviation from spherical symmetry for such rare peaks. This is very important to not only justify spherical symmetry assumption but also estimate the angular momentum of PBHs formed in radiation domination. Based on the perturbative analysis based on peak theory, De Luca, Desjacques, Franciolini, Malhotra and Riotto (2019) De Luca et al. (2019) concluded that the root mean square of the nondimensional Kerr parameter $a_{*}$ is of the order of $10^{-2}$, while Harada, Yoo, Kohri, Koga and Monobe (2021) estimated the root mean square of $a_{*}$ to be of the order of $10^{-3}$ and showed that a small fraction of PBHs of masses much smaller than the mass enclosed within the Hubble horizon, $M\ll M_{H}$, as a result of critical phenomena, can have much larger spins Harada et al. (2021).
@@
Figure 6: Contour map of the distribution functions of $(\chi_{\rm eff},q)$, $(\chi_{\rm eff},{\cal M})$ and $({\cal M},q)$, where $\chi_{\rm eff}$, ${\cal M}$ and $q$ are the effective spin parameter, the Chirp mass and the mass ratio of the binary PBHs, respectively. Taken from Ref. Koga et al. (2022).
@@
Based on this analysis, Koga, Harada, Tada, Yokoyama and Yoo (2022) Koga et al. (2022) calculated the distribution of the effective spin parameter $\chi_{\rm eff}$ of binary black holes, which is a well determined observable of the spins obtained from gravitational wave forms from inspiralling binaries. Figure 6 shows the distribution functions of the parameters $a_{*}$ and $M$ of the single PBHs and of $\chi_{\rm eff}$, the Chirp mass ${\cal M}$ and the mass ratio $q$ of the binary PBHs, respectively. Although there is a tendency for larger spins with the smaller masses because of the critical phenomena, the probabilities for large values of the spin parameters are strongly suppressed. See Ref. Koga et al. (2022) for details.
@@
In the above analyses, it was assumed that the power spectrum of curvature perturbation is sharp in the momentum space. Recently, it has been revealed that if this assumption is relaxed, the root mean square of $a_{*}$ can be slightly larger for some set of broad power spectra than the sharp one but should still be bounded by the value of the order of $10^{-3}$ Banerjee and Harada (in preparation).
@@ 4.3 Spins of primordial black holes formed in (early) matter domination
Since PBH formation will be enhanced in the (early) matter-dominated phase, it is very important to predict the spins of them. As we have seen before, nonspherical effects may become important. In fact, Harada, Yoo, Kohri and Nakao (2017) Harada et al. (2017) investigated the effects of angular momentum for PBH formation in this phase. Both the first-order and second-order effects can potentially play important roles. The first-order effect generates angular momentum through the nonsphericity of the region to collapse to a black hole, which is generally misaligned with a mode of linear perturbation. This is schematically illustrated in Fig. 7. The second-order does it through the coupling of two independent modes of linear perturbation, which is illustrated in Fig. 1 of Ref. Harada et al. (2017).
Figure 7: Schematic figures of the first-order effect to generate angular momentum. $V$ denotes the region that will collapse.
@@
where $a_{*}$ is the nondimensional spin parameter of the region to collapse and $\sigma_{H}$ is the standard deviation of $\delta_{H}$ at the the horizon entry. Although the nondimensional numerical factor of the order of the unity on the right-hand side should be determined yet, this implies that most of PBHs have spins $a_{*}=O(1)$ if $\sigma_{H}\sim 0.1$. The angular momentum effects will strongly suppress PBH formation if $\sigma_{H}$ is even much smaller because of the Kerr bound $|a_{*}|\leq 1$. These results have recently been updated based on peak theory Saito et al. (in preparation).
@@ 4.4 Nonspherical simulation of PBH formation
Yoo, Harada and Okawa (2020) Yoo et al. (2020) implemented 3D numerical simulation of nonspherical PBH formation in radiation domination based on numerical relativity for the first time. They prepared the long-wavelength solutions as initial data, for which initial nonsphericity is expected to be typically very small according to peak theory. However, to make the numerical results clearer, they put $\sim 10\%$ nonsphericity in the initial data, which is much larger than the values expected from peak theory for PBH formation. They found that even such large nonsphericity changed the threshold of PBH formation only by $\sim 1\%$. See Figs 1 and 3 of Ref. Yoo et al. (2020) for the initial density perturbation on the left panel and the apparent horizon formation in the course of gravitational collapse on the right panel. We can see that in the simulation with the near-threshold value a very small apparent horizon forms at the very central region. This implies that we need very high resolution near the centre. This problem was attacked with a rescaled radial coordinate Yoo et al. (2020).
Yoo (2024) Yoo (2024) implemented further numerical relativity simulations for the EOS $p=w\rho$ with $w=0.2$ and $1/3$ with much higher resolution and accuracy. For this purpose, not only the rescaled radial coordinate but also the multi-level mesh refinement scheme were adopted. In such simulations, the estimation of the spin is not straightforward. In this work, it was implemented as follows. For the Kerr black hole, we can write the spin parameter $a_{*}$ using the event horizon configuration through the following relation
@@ 5.1 Positive curvature region, type II perturbation and separate universe condition
In the expanding universe, we assume that the decaying modes of perturbation can be neglected, while the growing modes will lead to structure formation including the formation of PBHs. The growing mode of perturbation with positive density perturbation will undergo the maximum expansion to form a self-gravitating object. Such regions can locally be well modelled by a positive-curvature FLRW spacetime surrounded by a flat FLRW spacetime Carr and Hawking (1974).
There are two independent physical length scales to characterise the spatial geometry, the curvature radius and the size of the curved region. This consideration implies the separate universe configuration if the latter becomes as large as possible Carr and Hawking (1974) as schematically illustrated in Fig. 8. It should be noted that this figure does not show the sequence of time evolution but each configuration gives a set of initial data for each time development of perturbation. It is physically interesting to think a nearly separate universe configuration, while a totally separate universe has nothing to do with our observable universe. Kopp, Hofmann and Weller (2010) Kopp et al. (2011) classified the configuration of the spatial geometry into type I and type II as shown in Fig. 8. The marginal case is given by a 3-hemisphere, while smaller and larger 3-spheres are called types I and II, respectively. The marginal configuration gives the possible maximum amplitude of the density perturbation Kopp et al. (2011); Carr and Harada (2015), despite the initial claim that the separate universe condition does Carr and Hawking (1974). Since the curvature perturbation is larger for type II than for type I, type II configurations are much rarer than type I in a standard probability distribution function of the curvature perturbation, whereas type II can be dominant in a particular inflationary scenario Escrivà et al. (2023).
Figure 8: Classification of spatial configurations of overdense perturbations
@@ 5.2 Type II perturbation and its time development
Here we review the sets of initial data in terms of types I and II based on the recent work by Uehara, Escrivà, Harada, Saito and Yoo (2024) Uehara et al. (2024) for radiation domination. Here, the system is assumed to be spherically symmetric. They first constructed the long-wavelength solutions with choosing a function $\zeta(r)$. They chose the several choices for $\zeta$ but one of them is given by
@@
where $W(r)$ is an appropriate window function and the constant $\mu$ gives the amplitude of perturbation. The areal radius $R$ and the Shibata-Sasaki compaction function ${\cal C}_{\rm SS}$ are plotted as functions of $r$ in Figs. 1(b) and 2(a) of Uehara et al. (2024), respectively, where we can see that for $\mu\gtrsim 1.4$, $R$ is no longer a monotonic function of the rescaled radial coordinate $z$, which implies that there is a throat and that ${\cal C}_{\rm SS}(r)$ has two peaks with the value $1/2$ and a minimum in between. See Ref. Uehara et al. (2024) for the account for this peculiar behaviour of the compaction function. These features are essentially the same in matter domination as discussed in Ref. Kopp et al. (2011) using the LTB solution.
Then, the time development of thus constructed initial data was constructed with a standard numerical relativity scheme based on the Baumgarte-Shapiro-Shibata-Nakamura formalism but adjusted for the spherical formation of PBHs. Figure 3 of Uehara et al. (2024) summarises the evolution of the spacetime for the long-wavelength solutions generated by the curvature perturbation given by Eq. (23). For $\mu=0.5$, the amplitude of perturbation is so small that it cannot collapse but disperse away. For $\mu=1.2$, the the amplitude of perturbation is so large that it can collapse to a black hole. The growth of the density contrast as well as the rapid decrease to zero of the lapse function $\alpha$ indicate the continued collapse to a black hole. This suggests that the critical value of $\mu$ for the black hole formation is between $0.5$ and $1.2$. For $\mu=1.8$, the time evolution of the density contrast and the lapse function does not look so different at least in qualitatively from that for $\mu=1.2$.

### v3

@@ 4 Initial spins
If the black hole no hair conjecture holds in astrophysics, stationary black holes in vacuum should be well approximated by Kerr black holes which are characterised only by two parameters, the mass and the angular momentum. As we have already discussed, the mass of PBHs is approximately equal to that within the Hubble horizon at the time of formation. So, what determines the spin of PBHs? If black holes of masses from several to several tens of solar masses are observed, it would be very difficult to distinguish between PBHs and astrophysical black holes. Astrophysical black holes form in the final stage of evolution of massive stars. If we observe an isolated black hole, the only information of its own other than its mass is its spin. So, if the spins of PBHs are expected to be very different from those of astrophysical black holes, the observation of spin can be potentially decisive information to distinguish between the two populations. Here, we focus on the initial spins of PBHs. However, it should be noted that the accretion and merger history after their formation could also greatly affect their spins depending on the scenarios De Luca et al. (2020) and, hence, the theoretically estimated values for the initial spins should be compared with observation with great care.
@@ 4.1 Spins of primordial black holes formed in radiation domination
It was discussed that the formation process of PBHs in radiation domination is well approximated by spherically symmetric dynamics since Carr (1975) Carr (1975). Recent studies have revealed that this early argument is basically correct. Chiba and Yokoyama (2017) Chiba and Yokoyama (2017) gave an upper bound $\sim 0.4$ on the root mean square of the nondimensional Kerr parameter $a_{*}=a/M$, where $a$ and $M$ are the Kerr parameter and the mass of the black hole. He and Suyama (2019) He and Suyama (2019) discussed the effect of angular momentum on the threshold.
@@
Peak theory features more accurate quantitative analyses. For PBHs formed in radiation domination, the threshold density perturbation $\delta_{\rm th}$ is of the order of the unity, which is considerably larger than its standard deviation. In other words, PBHs can form only at very rare peaks. Peak theory predicts that there is only very small deviation from spherical symmetry for such rare peaks. This is very important to not only justify spherical symmetry assumption but also estimate the angular momentum of PBHs formed in radiation domination. Based on the perturbative analysis based on peak theory, De Luca, Desjacques, Franciolini, Malhotra and Riotto (2019) De Luca et al. (2019) concluded that the root mean square of the nondimensional Kerr parameter $a_{*}$ is of the order of $10^{-2}$, while Harada, Yoo, Kohri, Koga and Monobe (2021) estimated the root mean square of $a_{*}$ to be of the order of $10^{-3}$ and showed that a small fraction of PBHs of masses much smaller than the mass enclosed within the Hubble horizon, $M\ll M_{H}$, as a result of critical phenomena, can have much larger spins Harada et al. (2021).
@@
Figure 5: Contour maps of the distribution functions of $(M,a_{*})$ for isolated PBHs on the top panel, $M$ and $a_{*}$ are the mass and the nondimensional Kerr parameter, respectively, and $(\chi_{\rm eff},q)$, $(\chi_{\rm eff},{\cal M})$ and $({\cal M},q)$ for binary PBHs from left to right on the bottom panels, where $\chi_{\rm eff}$, ${\cal M}$ and $q$ are the effective spin parameter, the Chirp mass and the mass ratio of the binary PBHs, respectively. The definitions of $q$, ${\cal M}$ and $\chi_{\rm eff}$ are $q:=M_{2}/M_{1}$, ${\cal M}:=(M_{1}M_{2})^{3/5}/(M_{1}+M_{2})^{1/5}$ and $\chi_{\rm eff}:=(a_{*1}\cos\theta_{1}+qa_{*2}\cos\theta_{2})/(1+q)$, respectively, where $M_{i}$, $a_{*i}$ and $\theta_{i}$ ($i=1,2$) are individual masses, individual nondimensional Kerr parameters and the angles of individual spins with respect to the orbital angular momentum, respectively, with $M_{1}\geq M_{2}$ being assumed. Taken from Ref. Koga et al. (2022).
@@
Based on this analysis combined with peak theory, Koga, Harada, Tada, Yokoyama and Yoo (2022) Koga et al. (2022) calculated the distribution of the effective spin parameter $\chi_{\rm eff}$ of binary black holes, which is a well determined observable of the spins obtained from gravitational wave forms from inspiralling binaries. Figure 5 shows the distribution functions of the parameters $a_{*}$ and $M$ of isolated PBHs on the top panel and of the mass ratio $q:=M_{2}/M_{1}$, the Chirp mass ${\cal M}:=(M_{1}M_{2})^{3/5}/(M_{1}+M_{2})^{1/5}$ and $\chi_{\rm eff}:=(a_{*1}\cos\theta_{1}+qa_{*2}\cos\theta_{2})/(1+q)$ of binary PBHs on the bottom panels, where $M_{i}$, $a_{*i}$ and $\theta_{i}$ ($i=1,2$) are individual masses, individual nondimensional Kerr parameters and the angles of individual spins with respect to the orbital angular momentum, respectively, with $M_{1}\geq M_{2}$ being assumed. Although there is a tendency for larger spins with smaller masses because of the critical phenomena, the probabilities for large values of the spin parameters are strongly suppressed. See Ref. Koga et al. (2022) for details.
@@
In the above analyses, it was assumed that the power spectrum of curvature perturbation is nearly monochromatic. Recently, it has been revealed that if this assumption is relaxed, the root mean square of $a_{*}$ can be slightly larger for some set of broad power spectra but should still be bounded by the value of the order of $10^{-3}$ Banerjee and Harada (2024).
@@ 4.3 Spins of primordial black holes formed in (early) matter domination
Since PBH formation will be enhanced in the (early) matter-dominated phase, it is very important to predict the spins of them. As we have seen before, nonspherical effects may become important. In fact, Harada, Yoo, Kohri and Nakao (2017) Harada et al. (2017) investigated the effects of angular momentum for PBH formation in this phase. Both the first-order and second-order effects can potentially play important roles. The first-order effect generates angular momentum through the nonsphericity of the region to collapse to a black hole, which is generally misaligned with a mode wave number vector of the velocity perturbation. This is schematically illustrated in Fig. 6, where the ellipsoidal region that will collapse is misaligned with the wave number of the velocity perturbation denoted by the red arrows and, hence, it carries nonvanishing angular momentum. The secondary effect also gives angular momentum through the coupling of two independent modes of linear perturbation, which is illustrated in Fig. 1 of Ref. Harada et al. (2017).
Figure 6: Schematic illustration of the first-order effect to generate angular momentum. The blue region $V$ denotes the region that will collapse, which is approximated by an ellipsoid. The red arrows denote the velocity perturbation ${\bm{u}}$ with a wave number vector being parallel to the vertical direction. Since the ellipsoidal region $V$ is misaligned with the wave number ${\bm{k}}$ of the velocity perturbation, it carries nonvanishing angular momentum. The angular momentum can increase in time as the perturbation grows. This can also be understood through the effect of torque exerted on the boundary $\partial V$. See Ref. Harada et al. (2017) for details.
@@
where $a_{*}$ is the nondimensional spin parameter of the region to collapse and $\sigma_{H}$ is the standard deviation of $\delta_{H}$ at the the horizon entry. Although the nondimensional numerical factor of the order of the unity on the right-hand side should be determined yet, this implies that most of PBHs have spins $a_{*}=O(1)$ if $\sigma_{H}\sim 0.1$. The angular momentum effects will strongly suppress PBH formation if $\sigma_{H}$ is even much smaller because of the Kerr bound $|a_{*}|\leq 1$. These results have recently been updated based on peak theory Saito et al. (2024).
@@ 4.4 Nonspherical simulation of PBH formation
Yoo, Harada and Okawa (2020) Yoo et al. (2020) conducted 3D numerical simulation of nonspherical PBH formation in radiation domination based on numerical relativity for the first time. They prepared the long-wavelength solutions as initial data, for which initial nonsphericity is expected to be typically very small according to peak theory. However, to make the numerical results clearer, they put $\sim 10\%$ nonsphericity in the initial data, which is much larger than the values expected from peak theory for PBH formation. They found that even such large nonsphericity changed the threshold of PBH formation only by $\sim 1\%$. See Fig. 7 which is taken from Ref. Yoo et al. (2020) for the initial density perturbation on the left panel and the apparent horizon formation in the course of gravitational collapse on the right panel. We can see that in the simulation with the near-threshold value a very small apparent horizon forms at the very central region. This implies that we need very high resolution near the centre. This problem was attacked with a rescaled radial coordinate Yoo et al. (2020).
Refer to caption Refer to caption
Figure 7: The initial density perturbation on the left panel and the apparent horizon formation in the course of gravitational collapse on the right panel. We can see that an apparent horizon forms in the central region, which is very small compared to the size of the whole computational domain. It implies that to resolve the apparent horizon for the near-critical collapse requires very high resolution at least near the centre. Both taken from Ref. Yoo et al. (2020) with permission.
Yoo (2024) Yoo (2024) conducted further numerical relativity simulations for the EOS $p=w\rho$ with $w=0.2$ and $1/3$ with much higher resolution and accuracy. For this purpose, not only the rescaled radial coordinate but also the multi-level mesh refinement scheme were adopted. In such simulations, the estimation of the spin is not straightforward. In this work, it was implemented as follows. For the Kerr black hole, we can write the spin parameter $a_{*}$ using the event horizon configuration through the following relation
@@
As far as I am aware, no 3D numerical relativity simulations for nonspherical PBH formation in the (early) matter-dominated phase have been carried out using a fluid or dust description of the matter fields. In the scalar field approach to the early matter-dominated phase, De Jong, Aurrekoetxea, Lim and França (2023) de Jong et al. (2023) studied nonspherical PBH formation through 3D numerical relativity simulations and reported that the spin in the final stage of formation would be negligible. Although this result is highly suggestive, further systematic studies are anticipated in the near future.
@@ 5 Type II perturbation and type B PBH
5.1 Introduction to type II perturbation
In the FLRW solution, the geometry of the constant cosmological time $t$ spacelike hypersurface is given by a constant curvature space with a scale factor $a(t)$. The metric can be written in the familiar form:
$ds^{2}=-dt^{2}+a^{2}(t)\left[\frac{1}{1-K\tilde{r}^{2}}d\tilde{r}^{2}+\tilde{r}^{2}(d\theta^{2}+\sin^{2}\theta d\phi^{2})\right].$ (25)
The spatial curvature $K$ can be zero, positive or negative. For $K\neq 0$, the geometry has a finite curvature radius $1/\sqrt{|K|}$ up to the scale factor. The positive curvature solution generally has particular time evolution in $a(t)$, which begins to expand from big bang, reaches maximum expansion and collapses to big crunch, while for the zero and negative cases it just continues expanding from big bang to infinity. As is well known, for the positive-curvature FLRW spacetime, there is coordinate singularity at $\tilde{r}=1/\sqrt{K}$ in the line element (25). We can go beyond this singularity by introducing the radial coordinate $\chi$ in place of $\tilde{r}$ so that the line element can be written in the form:
$ds^{2}=-dt^{2}+a^{2}(t)[d\chi^{2}+\sin^{2}\chi(d\theta^{2}+\sin^{2}\theta d\phi^{2})],$ (26)
where $K$ is normalised to $1$ using the freedom in rescaling $a(t)$. We can see that the coordinate singularity which was at $r=1$ in the $(t,r)$ coordinates is now resolved at $\chi=\pi/2$ in the $(t,\chi)$ coordinates. Roughly speaking, a type II perturbation is the perturbation that covers the surface $\chi=\pi/2$, while a type I does not.
In the perturbation theory in the flat FLRW solution, there appear growing and decaying modes. We can usually neglect decaying modes in the early Universe if we are interested in subsequent structure formation including PBH formation. Then, we can expect that the regions with positive and negative density perturbations can be described by locally positive and negative curvature FLRW solutions, respectively. In fact, this is really true in the long-wavelength solutions because Eq. (14), which is the metric in the long-wavelength limit, can be rewritten in the following form Polnarev and Musco (2007); Harada et al. (2015):
$ds^{2}=-dt^{2}+a^{2}(t)\left[\frac{1}{1-K(\tilde{r})\tilde{r}^{2}}d\tilde{r}^{2}+\tilde{r}^{2}(d\theta^{2}+\sin^{2}\theta d\phi^{2})\right]$ (27)
in a new radial coordinate $\tilde{r}$. Since type II perturbations need larger amplitudes of curvature perturbation, they are much rarer than type I in a standard probability distribution function of the curvature perturbation, whereas type II can be dominant in a particular inflationary scenario with large nonGaussianties Escrivà et al. (2023).
@@ 5.2 Positive curvature region, type II perturbation and separate universe condition
As we have discussed, the region with a positive density perturbation can locally be well described by the positive-curvature FLRW solution at least in the long-wavelength limit. For understanding the geometry and the dynamics of such a system, we introduce a toy model, which consists of a positive-curvature FLRW region surrounded by a flat FLRW region Kopp et al. (2011); Harada et al. (2013). We call this model the 3-zone model. Figure 8 gives its schematic figure. The line elements in regions I and III are written in the following forms, respectively,
$ds^{2}=-dt^{2}+a^{2}_{\rm I}(t)\left[d\chi^{2}+\sin^{2}\chi(d\theta^{2}+\sin^{2}\theta d\phi^{2})\right]$ (28)
for $0<\chi<\chi_{a}$ and
$ds^{2}=-dt^{2}+a^{2}_{\rm III}(t)\left[dr^{2}+r^{2}(d\theta^{2}+\sin^{2}\theta d\phi^{2})\right]$ (29)
for $r_{b}<r$, where $a_{\rm I}(t)$ and $a_{\rm III}(t)$ are the scale factors of the positive-curvature and flat FLRW solutions, while that in region II cannot be written in a simple form. Although this toy model is not only unrealistic but also difficult to justify as the whole evolution of the spacetime unless we make strong assumptions on region II, it is still valid as the long-wavelength solution and useful to understand the concept of the nonlinear perturbation. In this model, we have two independent physical length scales of region I, the curvature radius, which is normalised to $1$ in the above metric, and the size of the region, which is given by $\chi_{a}$. As $\chi_{a}$ is increased from 00 to $\pi/2$, the comoving areal radius of region I, which is given by $\sin\chi_{a}$, monotonically increases from 00 to $1$. However, as $\chi_{a}$ is further increased from $\pi/2$ to $\pi$, the comoving areal radius turns to decrease from $1$ to 00. This consideration implies the separate universe configuration for which $\chi_{a}=\pi$, which is notified in Ref. Carr and Hawking (1974). Kopp, Hofmann and Weller (2010) Kopp et al. (2011) classified the configurations with $20<\chi_{a}<\pi/$ 2 and with $\pi/2<\chi_{a}<\pi$ into types I and II, respectively, while the marginal case is given by a 3-hemisphere $\chi_{a}=\pi/2$, where we neglect the contribution of region II.
Figure 8: The 3-zone model of the positive density perturbation. Regions I and III are described by the closed and flat FLRW solutions, while region II is an underdense matching layer. The spheres of $\chi=\chi_{a}$ and $r=r_{b}$ give the outer edge of region I and the inner edge of region III, respectively. See Refs. Kopp et al. (2011); Harada et al. (2013) for more details.
Apart from this toy model, it is still true that there should be two independent physical length scales to characterise the spatial geometry, the curvature radius and the size of the curved region as schematically illustrated in Fig. 9. It should be noted that this figure does not show the sequence of time evolution but each configuration gives a set of initial data for each time development of perturbation. It is physically interesting to think a nearly separate universe configuration, while a totally separate universe has nothing to do with our observable Universe.
Figure 9: Classification of spatial configurations of overdense perturbations. The type II configuration is with the throat structure in the spatial geometry, while the type I is not. Note that the sequence does not correspond to the time evolution but to different initial curvature perturbations.
@@ 5.3 Type II perturbation and its time development
Here we review the sets of initial data in terms of types I and II based on the recent work by Uehara, Escrivà, Harada, Saito and Yoo (2024) Uehara et al. (2024) for radiation domination. Here, the system is assumed to be spherically symmetric. They first constructed the long-wavelength solutions with choosing a function $\zeta(r)$. One of the functional forms they tried for $\zeta(r)$ is given by
@@
where the constant $\mu$ gives the amplitude of perturbation and $W(r)$ is an appropriate window function chosen so that the Gaussian tail can be eliminated and $\zeta$ can be smoothly matched to 00 at the outer boundary of the numerical domain. The areal radius $R$ and the Shibata-Sasaki compaction function ${\cal C}_{\rm SS}$ are plotted as functions of $r$ in Figs. 1(b) and 2(a) of Uehara et al. (2024), respectively, where we can see that for $\mu\gtrsim 1.4$, $R$ is no longer a monotonic function of the radial coordinate $r$, which implies that there is a throat and that ${\cal C}_{\rm SS}(r)$ has two peaks with the value of $1/2$ and a minimum in between. See Ref. Harada et al. (2024); Uehara et al. (2024) for the account for this peculiar behaviour of the compaction function. These features are essentially the same in matter domination as discussed in Ref. Kopp et al. (2011) using the LTB solution.
Then, the time development of those initial data was constructed with a standard numerical relativity scheme based on the Baumgarte-Shapiro-Shibata-Nakamura formalism but adjusted for the spherical formation of PBHs. Figure 3 of Uehara et al. (2024) summarises the evolution of the spacetime for the long-wavelength solutions generated by the curvature perturbation given by Eq. (30). For $\mu=0.5$, the amplitude of perturbation is so small that it cannot collapse but disperses away. For $\mu=1.2$ and $1.8$, the the amplitude of perturbation is so large that it can collapse to a black hole. This suggests that the critical value of $\mu$ for the black hole formation is between $0.5$ and $1.2$.